OpenStudy (cicilybailey14):

PLEASE HELP!!! Molly paid $1,055.60 for a used car. If the price paid includes a 9% discount, which of the following equations can be used to determine the price of the car before the discount? (Let x represent the cost of the car and y represent the total cost before the discount.)

OpenStudy (triciaal):

original y = 100% the whole 9% discount x = (100-9)= 91%Y

OpenStudy (triciaal):

x = 1055.60 1055.60 = 0.91 Y divide each side by 0.91 y =

OpenStudy (triciaal):

which of the following equations I don't have the choices but you should be able to see which one from the info above

OpenStudy (cicilybailey14):

A. y = 1.09x B. y = 1.9x C. y = 0.91x D. y = x + 9x